# An open framework for archival, reproducible, and transparent science

### v1

@@ 3.4 Trusted timestamping
Trusted timestamping (RFC3161/RFC5816) is a technique that provides secure tracking of creation and modification dates for any digital content including data and code [38]. Since data and code have different life cycles, the trusted timestamping process differs between them.
Data should be timestamped during the collection process. This can be achieved using the scripts available in the Trusted Timestamping Framework for Scientific Research [2] or manually with the associated web tool [1]. The framework generates a timestamps.json file which should be included at the top-level directory where timestamped data is present. An example timestamped dataset is provided for reference [39].
@@ 3.5 Documentation
From the standpoint of replicability, clear methodology defining the data collection process should also be present. The fine details of this process, if included in research articles, are often left to supplementary sections and not usually accompanied by data collection pipelines. Artifact documentation should include technical details around data collection such as experimental setup, materials (including specific hardware) needed, and any related code.
@@ 3.6 Configuration and file structure
An environment variables file named config.env must be present at the top level folder of the deposition or under a folder named arts/ to be compatible with the ARTS open framework. This file defines where code, data, and research outputs are located, locations for other required files, and any other parameters that may be passed to analysis or data collection code. All other files and folders, while recommended to include when relevant, are optional and may be omitted.
@@
Researchers are responsible for using these environment variables in their code so that changing them has an affect on analysis reruns. Omitting an environment variable or setting it to an empty string (e.g., ARTS_CODE_PATH= ) indicates that corresponding file or folder is not present in the deposition. In some circumstances, researchers may also want to provide a PURL or other resolvable value in path variable. For example, if new features are being derived from an existing raw dataset, the ARTS_RAW_DATA_PATH may be set to an external DOI or PURL where the data is accessible, so long as the setup and run commands support using an external URL.
@@
Figure 3: Minimal example file structure adhering to the ARTS open framework.
The presence of a config.env greatly simplifies the reproduction process by allowing validators to rerun analyses and recreate the output/ directory with the one-line ARTS_RUN_CMD. Replication is also largely simplified for results that are reproducible through the ARTS open framework. After running a new experiment to collect independent data, the value for ARTS_RAW_DATA_PATH can be updated to run analyses on these data and complete the replication. Incorporating top-level environment variables during the analysis and data collection processes is also a recommended best practice as it provides a unified interface for parameter tuning and makes code more accessible.
@@ 3.7 Licensing
As for data and other content such as documentation, Creative Commons (CC) provides three licenses that are widely used and recommended. Open Data Commons (ODC) also provides a set of similar open licenses that are aimed specifically at data, but may not be suitable for other content [41]. The most recent versions of the CC licenses are recommended over ODC licenses as they provide broader coverage and permissions [9]. For placing data in the public domain without restriction, the CC0 license should be used, and is required for work directly generated by US federal agencies and staff. If attribution to the original authors is required during redistribution, the CC BY license is suitable. And if an attribution and share-alike clause is desired, there is the CC BY-SA license [10]. Unfortunately, none of these licenses provide terms for data provenance beyond attribution. For example, even a dataset released under the CC BY-SA license can be downloaded, slightly modified, and reuploaded as a purported exact copy of the original so long as it is released with attribution information and a copy of the license. Therefore, it is important to release attestable provenance information along with datasets as described in the trusted timestamping section.
@@
Figure 4: Open licenses apply to software and data and can require attribution, or openness of derivative work. Permissive licenses assign copyright, but give users broad usage and distribution rights whereas works placed in the public domain release copyright from the original author. Creative common licenses are popular for data and content, but open-source code is more often licensed under one of the other listed licenses.
@@
When choosing a license, there are a few important considerations to make [14]. There may be requirements set by government or foundational funding sources requiring research outputs be released under specific terms [53, 16, 26]. Specific archives may also have license requirements—for example, Dryad places all depositions under a CC0 license. After fulfilling any obligations, the intended audience of the code and data should be considered. Results that are intended to be built upon in a research setting may be best served under a share-alike license to encourage future openness. To maximize reach across commercial settings, placing code under permissive licenses and data in the public domain is recommended. Conversely, a noncommercial license may be used if the goal is to discourage commercial use. Dual licensing may also be an option for authors that wish to retain the ability to license their work commercially under stricter terms [52]. For research outputs, it is generally best to use a license that requires attribution at a minimum to preserve the historical record. An overview of recommended open licenses is provided in Figure 4.
4 Recommendations
@@
The following recommendations apply to specific groups of stakeholders involved with running, archiving, sharing, and funding research.
4.1 Researchers
@@
Researchers should fully document their workflows if they do not do so already. This includes spelling out all versions used, including packages and libraries used in code, programming languages, and operating systems tested on. Shell scripts that define and run the workflow should also be provided. Finally, providing a container image that bundles all aspects of the experiment is necessary in order to create a community around reproducing results and sharing scientific knowledge as broadly as possible. Implementing the ARTS open framework is a good way to achieve this and for well-designed computational studies, should primarily require specifying a config.env file and modifying associated top-level scripts.
4.2 Archivists
@@
Those who administer deposition archives should make sure that they have at least the minimal feature set described by Figure 2. Except for hosting container registry endpoints, a few archives are close to being full-featured in this regard. Zenodo and the Dataverse Project should expose file structure in URLs to support git repo cloning. The Internet Archive could be a suitable archive for research projects if it added support for immutable releases and temporary private uploads for the purpose of peer review. Stanford’s Digital Repository would also support all aspects of the standard if it opened up submissions to external collaborators. Building in useful provenance tools like trusted timestamping also moves some of the work off of researchers and increases the reputability of the archive. All major archiving platforms should support pulls for OCI containers.
4.3 Publishers
@@
Publishers and those involved with the scientific review process should also require that the research they are distributing conforms to reproducibility best practices. The peer review process for submissions that include code should involve having an independent recreation of relevant figures from referenced data. A number of Nature journals have taken a step in this direction by integrating and encouraging the use of Code Ocean for reviews [34, 33]. Publishers, especially those already partnered with archive services, should also be aware of the opportunity to implement automated ”clearing house” services that run the deposited container to validate figure generation from data for ARTS-compatible submissions. Encouraging third-party figure recreation in this manner sets a high bar for computational reproducibility.
4.4 Funding agencies and policy makers
@@
Many US funding agencies are taking steps to ensure that data generated from sponsored awards be made public at the time of publication [36]. While this is an excellent first step, simply having the data released does not mean the data is usable or accessible. Code and the computational environment needed to generate the outputs of research could also be required to be made public in the same manner, and this would significantly increase the accessibility of research works. The ARTS framework would be an example framework compliant with both the letter and the spirit of such requirements.
5 Example Implementation: Watch Calibration
@@
A 2813 automatic mechanical movement was recorded with a laptop by placing the movement and case on the interior of the laptop above the headphone jack where the microphone is located. Other watch movements and recording equipment may be used, but when recording with a laptop it’s important to place the watch directly on top of the laptop microphone and record in a quiet environment. Audio data was recorded at 44100 Hz using the sounddevice Python package. The experiment code also offers the ability to generate a dataset using the librosa.clicks method [31]. Derivative data in the form of audio windows containing each clock tick and their respective peak onset time were output to the ARTS_DERIV_DATA_PATH folder.
Figure 5: Comparison of drift values across each dataset recorded (W241130; 250227) and generated (250409; 250310_noisy) for all possible analysis parameters. Drifts are fairly consistent for recordings on the left with high SNR. Data with more noise maybe need filtering for a reasonable drift calculation (250410_noisy) or filtering may result in additional noise (250227).
@@
The container containing the analysis can be run via Podman or Docker using the ./run.sh script available in the deposition [11]. The run script provides a command to regenerate figures (./run.sh generate-figures), launch a jupyter notebook (./run.sh jupyter), or launch an ipython interactive shell (./run.sh ipython). The recording process can also be rerun using the collect_data.sh script to collect new data using a system microphone or generate new data using librosa. The container image may also be regenerated with ./run.sh save-image.
6 Limitations and Future Work
6.1 Limitations
The ARTS open framework is not a silver bullet for research reproducibility and replicability, but simply a set of flexible best practices. It is not a replacement for careful data collection, statistical rigor, or clear explanation of methods. The framework is dependent on a number of other technologies, notably: git (although it may be extensible to other version control systems), the OCI’s container specifications, and chosen archive services. It also does not yet adequately address workflows that work with large raw data on the order of terabytes or more per experiment or those that involve specific or proprietary hardware.
And although the ARTS open framework offers wide support across compute environments, steps to perform GPU workflows or hybrid workflows that span local and cloud compute clusters are not prescribed in this manuscript or accompanying code. Further, container images generated on x86 host systems will only reliably run on other x86 systems and container support across various computing architectures may be limited [46].
It should also be noted that container images for scientific computing tend to be quite large (on the order of 1-10 GB depending on the included dependencies). Archives serving these files can take advantage of layer caching—the process of sharing reusable blocks of containers that share dependencies—to greatly reduce the total storage needed to host container images.
6.2 Future Work
ARTS is a generic framework, but to make life easy for researchers in their respective fields, it’s important to have opinionated frameworks. One signal processing example experiment using the framework is provided at the time of writing, but for the ARTS open framework to be most useful, there must be further examples created adhering to the framework that detail specific data collection pipelines and shared environments for analyses using programming languages besides Python, such as R and MATLAB. Clear examples on how to archive large data on the order of terabytes with trusted timestamping and version control need also be added.
7 Conclusion
Science, especially that which is funded by the public, is an asset that should be accessible to the public including other researchers. Although new standards and guidelines make the outputs of scientific endeavors available, they are often not easily usable, even to other experts in the field. This large burden of making results reproducible currently falls on researchers and highlights the importance of tools that are compatible with already-used workflows.
To bridge this knowledge gap requires education on both sides. Researchers and interested end users need to learn at least the high-level DevOps principles necessary to output their work in an accessible format and run those outputs. Fluency with these techniques are often left out of university curricula and treated as professional skills, but they are now core to modern science.
Given that no current research artifact archive meets all recommendations of the ARTS open framework, archive services are encouraged to preserve deposition file structure in URLs and move towards supporting container registries that house the container images used to validate ARTS-compatible research outputs. Implementing these features will allow for automatic recreation of figures and validation of computational results.
By providing a framework for reproducible workflows, the hope is to enable other researchers to further increase the replicability of their work and encourage the focus of replication to be on sound scientific methods for robust data collection and not interpretation of analysis.

### v2

@@ 3.5 Trusted timestamping
Trusted timestamping (RFC3161/RFC5816) is a technique that provides secure tracking of provenance information (i.e., creation and modification dates) for any digital content including data and code [36]. Since data and code have different life cycles, the trusted timestamping process differs between them.
Data should be timestamped during the collection process. A reference implementation is available using the scripts in the Trusted Timestamping Framework for Scientific Research [56] or manually with the associated web tool [57]. The framework generates a timestamps.json file which should be included at the top-level directory where timestamped data is present. An example timestamped dataset is provided for reference [58].
@@ 3.6 Documentation
From the standpoint of replicability, clear methodology defining the data collection process should also be present. The fine details of this process, if included in research articles, are often left to supplementary sections and not usually accompanied by data collection pipelines. Artifact documentation should include technical details around data collection such as experimental setup, materials (including specific hardware) needed, and any related code. To the extent possible, all code should also be well documented in accordance with the respective programming language guidelines.
@@ 3.7 Configuration and file structure
An environment variables file named config.env or .config.env must be present at the top level folder of the deposition or under a folder named arts/ to be compatible with the ARTS open framework. This file defines where code, data, and research outputs are located, locations for other required files, and any other parameters that may be passed to analysis or data collection code. All other files and folders, while recommended to include when relevant, are optional and may be omitted.
@@
Researchers are responsible for using these environment variables in their code so that changing them has an affect on analysis reruns. In the ideal case, where this framework is incorporated early on in the research lifecycle, these path variables point to local resources and can later be updated to reflect archive deposition compatible paths. If the downstream code uses these variables for all path definitions, then the code will successfully port to any environment or archive that the deposition may be moved to. Omitting an environment variable or setting it to an empty string (e.g., ARTS_CODE_PATH= ) indicates that the corresponding file or folder is not present in the deposition. In some circumstances, researchers may also want to provide a PURL or other resolvable value in a path variable. For example, if new features are being derived from an existing raw dataset, the ARTS_RAW_DATA_PATH may be set to an external DOI or PURL where the data is accessible, so long as the setup and run commands support using an external URL.
@@
Figure 3: Minimal example file structure compatible with the ARTS open framework.
The presence of a config.env greatly simplifies the reproduction process by allowing validators to rerun analyses and recreate the output/ directory with the one-line ARTS_RUN_CMD. Replication is also largely simplified for results that are reproducible through the ARTS open framework. After running a new experiment to collect independent data, the value for ARTS_RAW_DATA_PATH can be updated to run analyses on these data and complete the replication. Incorporating top-level environment variables during the data collection and analysis development processes is also a recommended best practice as it provides a unified interface for parameter tuning and makes code more accessible. For example, when archiving sensitive data such as PHI or trade secrets, access controls such as authentication are necessary, but can be accomplished via an API key supplied as an environment variable.
@@ 3.8 Licensing
As for data and other content such as documentation, Creative Commons (CC) provides three licenses that are widely used and recommended. Open Data Commons (ODC) also provides a set of similar open licenses that are aimed specifically at data, but may not be suitable for other content [63]. The most recent versions of the CC licenses are recommended over ODC licenses as they provide broader coverage and permissions [64]. For placing data in the public domain without restriction, the CC0 license should be used, and is required for work directly generated by, or involving as authors, US federal agencies and staff. If attribution to the original authors is required during redistribution, the CC BY license is suitable. And if an attribution and share-alike clause is desired, there is the CC BY-SA license [65]. Unfortunately, none of these licenses provide terms for data provenance beyond attribution. For example, even a dataset released under the CC BY-SA license can be downloaded, slightly modified, and reuploaded as a purported exact copy of the original so long as it is released with attribution information and a copy of the license. Therefore, it is important to release attestable provenance information along with datasets as described in the trusted timestamping section.
@@
Figure 4: Open licenses apply to software and data and can require attribution, or openness of derivative work. Permissive licenses assign copyright, but give users broad usage and distribution rights whereas works placed in the public domain release copyright from the original author. Creative Commons licenses are popular for data and content, but open-source code is more often licensed under one of the other listed licenses.
@@
When choosing a license, there are a few important considerations to make [62]. There may be requirements set by government or foundational funding sources requiring research outputs be released under specific terms [66, 67, 68]. Specific archives may also have license requirements—for example, Dryad places all depositions under a CC0 license. After fulfilling any obligations, the intended audience of the code and data should be considered. Results that are intended to be built upon in a research setting may be best served under a share-alike license to encourage future openness. To maximize reach across commercial settings, placing code under permissive licenses and data in the public domain is recommended. Conversely, a noncommercial license may be used if the goal is to discourage commercial use outside of paid license agreements. Dual licensing may also be an option for authors that wish to retain the ability to license their work commercially under stricter terms [69]. For research outputs, it is generally best to use a license that requires attribution at a minimum to preserve the historical record. An overview of recommended open licenses is provided in Figure 4.
4 Discussion
4.1 Recommendations
@@
The following recommendations apply to specific groups of stakeholders involved with practicing, archiving, sharing, and funding research.
4.1.1 Researchers
@@
Researchers should work towards sharing the computing environments that they perform their research in. They should fully document their workflows which includes spelling out all dependency versions used, including packages and libraries used in code, programming languages, and operating systems tested on. Shell scripts that define and run the workflow should also be provided. Finally, providing a container image that bundles all aspects of the experiment is necessary in order to create a community around reproducing results and sharing scientific knowledge as broadly as possible. Implementing the ARTS open framework is a good way to achieve this and for well-designed computational studies, should primarily require specifying a config.env file and modifying associated top-level scripts.
4.1.2 Archivists
@@
Those who administer deposition archives should make sure that they have at least the minimal feature set described by Figure 2. Except for hosting container registry endpoints, a few archives are close to being full-featured in this regard. Zenodo and the Dataverse Project should expose file structure in URLs to support git repository cloning. The Internet Archive could be a suitable archive for research projects if it added support for immutable releases and temporary private uploads for the purpose of peer review. Stanford’s Digital Repository would also support almost all aspects of the standard if it opened up submissions to external collaborators. Incorporating useful provenance tools like trusted timestamping also moves some of the work off of researchers and increases the reputability of the archive. All major archiving platforms should support pulls for OCI container images.
4.1.3 Publishers
@@
Publishers and those involved with the scientific review process should require that the research they are distributing conforms to reproducibility best practices. The peer review process for submissions that include code should involve having an independent recreation of relevant figures from referenced data. A number of Nature Publishing Group journals have taken a step in this direction by integrating and encouraging the use of Code Ocean for reviews [70, 71]. Publishers, especially those already partnered with archive services, should also be aware of the opportunity to implement automated “clearing house” services that run the deposited container to validate figure generation from data for ARTS-compatible submissions. Encouraging third-party figure recreation in this manner sets a high bar for computational reproducibility.
4.1.4 Funding agencies and policy makers
@@
Many US funding agencies are taking steps to ensure that data generated from sponsored awards be made public at the time of publication [72]. While this is an excellent first step, simply having the data released does not mean the data is usable or accessible. Ideally, code and the computational environment needed to generate the outputs of research should also be required to be made public in the same manner, and this would significantly increase the accessibility of research works. The ARTS framework would be an example framework compliant with both the letter and the spirit of such requirements.
4.2 Limitations and Future Work
4.2.1 Limitations
The ARTS open framework is not a silver bullet for research reproducibility and replicability, but simply a set of flexible best practices. It is not a replacement for careful data collection, statistical rigor, or clear explanation of methods. The framework is dependent on a number of other technologies, notably: git (although it may be extensible to other version control systems), the OCI’s container specifications, and chosen archive services. It also does not yet adequately address workflows that work with large raw data on the order of tens of terabytes or more per experiment or those that involve specific or proprietary hardware.
And although the ARTS open framework offers wide support across compute environments, steps to perform GPU workflows or hybrid workflows that span local and cloud compute clusters are not prescribed in this writing or accompanying code. Further, container images generated on one host architecture may not always reliably run on other host architectures [73].
It should also be noted that container images for scientific computing tend to be quite large (potentially up to tens of gigabytes depending on the included dependencies). Archives serving these files can take advantage of layer caching—the process of sharing reusable blocks of containers that share dependencies—to greatly reduce the total storage needed to host container images.
4.2.2 Future Work
ARTS is a generic framework, but to make adoption easier and more accessible for researchers in their respective fields, it’s important to have opinionated examples. One signal processing example experiment using the framework is provided at the time of writing, but for the ARTS open framework to be most useful, there must be further examples created that are compatible with the framework that detail specific data collection pipelines and shared environments for analyses using programming languages besides Python, such as R and Julia. Examples of and documentation for ARTS-compatible depositions and the associated upload process across different archives including depositing within a cloud storage bucket would also make the framework easier to use. Clear examples on how to archive large data on the order of terabytes with trusted timestamping and version control need also be added.
5 Methods
5.1 Example Implementation: Watch Calibration
@@
A 2813 automatic mechanical movement was recorded with a laptop by placing the movement and case on the interior of the laptop above the headphone jack where the microphone is located. Other watch movements and recording equipment may be used, but when recording with a laptop the best results are obtained with the watch directly on top of the laptop microphone and a quiet recording environment. Audio data was recorded at 44100 Hz using the sounddevice Python package. The experiment code also offers the ability to generate a dataset using the librosa.clicks method [76]. Derivative data in the form of audio windows containing each clock tick and their respective peak onset time were output to the ARTS_DERIV_DATA_PATH folder.
Figure 5: Comparison of drift values across each dataset recorded (W241130; 250227) and generated (250409; 250310_noisy) for all possible analysis parameters. Drifts are fairly consistent for recordings on the left with high SNR. Data with more noise may need filtering for a reasonable drift calculation (250410_noisy) or filtering may result in additional noise (250227).
6
@@
The analysis can be run in its container environment via Podman or Docker using the ./run.sh script available in the deposition [77]. The run script provides a command to regenerate figures (./run.sh generate-figures), launch a jupyter notebook (./run.sh jupyter), or launch an ipython interactive shell (./run.sh ipython). The recording process can also be rerun using the collect_data.sh script to collect new data using a system microphone or generate new data using librosa. The container image may also be regenerated with ./run.sh save-image.
6 Data Availability
All raw data used in the example experiment is available in the SDR (https://doi.org/10.25740/vs897sz1847, https://purl.stanford.edu/vs897sz1847), on Zenodo (https://zenodo.org/records/15232081), or in the Internet Archive (https://purl.archive.org/watch-calibration).
7 Code Availability
All code used in the example experiment is available in the SDR (https://doi.org/10.25740/vs897sz1847, https://purl.stanford.edu/vs897sz1847), on Zenodo (https://zenodo.org/records/15232081), or on GitHub (https://github.com/bil/watch-calibration).
The full git repository can be cloned with git clone https://stacks.stanford.edu/file/vs897sz1847/watch-calibration/watch-calibration.git, but data must be downloaded separately. Note that this link is not a PURL, but the root URL may be found by navigating to the SDR deposition DOI/PURL and copying the download link for an individual file.